Find the slope and the y-intercept of the line. Write fractional answers in lowest terms.
3y+7x+5=9+7x
The slope is _____
The y-intercept is (0,______)

Answer by checkley77(12844)

3y+7x+5=9+7x
3y=7x-7x+9-5
3y=4
y=4/3
the slope (m)=0
the y intercept(b)=4/3